__ + __ + __ = 30 
Using 1,3,5,7,9,11,13,15 only
reYou can repeat the numbers too!

Grade:Select Grade

13 Answers

Anubhab Das
9 years ago
This is not possible. Three odd numbers can’t sum up to form an even number. :) 
Moksh
7 years ago
15+13+(1+1)=3015 and 13 is oddand we can repeat no. so 1 repeat twice in a braket to get 30................................
Rohit Goswami
6 years ago
11+13+3!=303!=3×2×1__________₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹++-₹&*^_|$^++^-*&^₹₹&&&%|$^*&$^*&%|*^$#$&&$$&&%|&*%#×+_₹^(^₹$#^&*^#%&*^$^**%%8*%₹*+:₹₹&((^₹#$$$₹₹₹₹₹₹₹₹₹_______₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹-₹$$$$$$%$$$$$₹₹_
Rohit Goswami
6 years ago
11+13+3!=30````` 3!=3×2×1=6,+~×~×××-_+××+++××××××+++₹₹₹_₹₹₹₹₹₹₹₹₹₹₹______::::₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹_______₹_₹__₹₹₹₹_₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹₹
Khaidem Kajalkumar Meitei
6 years ago
11 + 3! (1*2*3, which equals 6) + 13 = 30. Using Factorial Equations, it works. But, adding only odd numbers do not get any Even Number Like 30 for its Answer.
Niraj kumar saini
6 years ago
11-1+13-3+15-5=30 Correct ans this question Its right answer.................................................................. Totally odd manually
Avron Vernon Souto
6 years ago
 
15+13+(1+1)=30 , 15 and 13 are odd and i repeated 1 in brackets …......................................................................................................
Aayuska Shakya
6 years ago
(3+7)+(9+1)+(15-5)=30We can use subtraction method inside the brackets so we will obtain 30 as answer ✔
Dakshinamurti pandey
6 years ago
(15-1)+3+13=30 we can use subtraction method in this answer. We can add minus in the braket to bring 30. Other wise we cannot find any answer.
Amit
6 years ago
(15-1)+13+3=30Vhsigdphudtkydvgjxgjzirsurfhsgvdkfrnxgshgvdfkxhtl glcjfufjc vbg ny fw gshuegvs euhapgeuggdoqfdigsoja duhecduehdykdylpdhd f
Sharlee
6 years ago
9+11+(15-5)=30 we can use both subtraction and addition method in the brackets to obtain the answer 30
Ruturaj Hajare
6 years ago
15+15+(-30)=0....... We can take 30 at left side... No rules are broken.. And we are repeated 15 twise..
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
We must note that the sum of 3 odd numbers can never be even.
But given if each place can have more than one number then it might be possible.
For example-
(3+7)+(9+1)+(15-5)=30
9+11+(15-5)=30
(15-1)+3+13=30
And many more....
 
Thanks and regards,
Kushagra